# Thresholded Lasso for high dimensional variable selection

### v2

@@ 3 Proof sketch for the main result
Suppose all conditions in Theorem 2.4 hold. Then
Then by Lemma 3.1 and (53), we have
$\;\text{where}\;\ell(s_{0})$ ::=:= $\frac{\theta_{s_{0},2s_{0}}}{\sqrt{\Lambda_{\min}(2s_{0})}}\wedge\sqrt{\Lambda_{\max}(s_{0})}.$ (54)
Lemma 3.4.
Let $T_{1}$ be the $s_{0}$ largest positions of $h$ outside of $T_{0}$, where $T_{0}$ denotes the largest $s_{0}$ coordinates of $\beta$ in absolute values. Denote by $T_{01}=T_{0}\cup T_{1}$. Under the settings of Theorem 2.4, we have under event ${\mathcal{T}}_{a}$, for $h=\hat{\beta}-\beta_{0}$,
$\text{\bf Case 1}:\;\left\lVert h\right\rVert_{2}$$\leq$ $2\left\lVert h_{T_{01}}\right\rVert_{2}\leq 2D_{0,a}\lambda\sigma\sqrt{s_{0}}\;\text{where}$
$D_{0,a}$$\leq$ $\left\{\sqrt{2}\big{(}\sqrt{\Lambda_{\max}(s-s_{0})}K(s_{0},3)+3d_{0}K^{2}(s_{0},3)\big{)}\right\};$ (55)
$\text{\bf Case 2}:\;\left\lVert h\right\rVert_{2}$$\leq$ $\sqrt{5}\left\lVert h_{T_{01}}\right\rVert_{2}\leq\sqrt{10}K^{2}(s_{0},4)d_{0}\lambda\sigma\sqrt{s_{0}}$
$\text{\bf Case 3}:\;\left\lVert h_{T_{01}}\right\rVert_{2}$$\leq$ $\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\big{(}\left\lVert Xh\right\rVert_{n}+\ell(s_{0})\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}\big{)}\leq D\lambda\sigma\sqrt{s_{0}},$
$\text{where}\;D$ ::=:= $\frac{\sqrt{\Lambda_{\max}(s-s_{0})}}{\sqrt{\Lambda_{\min}(2s_{0})}}\big{(}1+\frac{3\ell(s_{0})}{d_{0}}\sqrt{\Lambda_{\max}(s-s_{0})},\big{)}$ (56)
and $\ell(s_{0})$ is as defined in (54). Moreover, under event ${\mathcal{T}}_{a}$, we have for $D_{0}$ in (2.1) and $D_{1}$ in (26): $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq D_{1}\lambda\sigma s_{0}$,
@@
$\left\lVert h\right\rVert_{2}^{2}$ $=$ $\left\lVert h_{T_{01}}\right\rVert_{2}^{2}+\left\lVert h_{T_{01}^{c}}\right\rVert_{2}^{2}\leq(D_{0}^{2}+D_{1}^{2})\lambda^{2}\sigma^{2}s_{0},\text{ and }\;$
$\left\lVert\hat{\beta}-\beta\right\rVert_{2}$$\leq$ $\left\lVert h\right\rVert_{2}+\lambda\sigma\sqrt{s_{0}}\leq[\sqrt{D_{0}^{2}+D_{1}^{2}}+1]\lambda\sigma\sqrt{s_{0}}.$
@@
We demonstrate the tightness of these bounds in Fig. 2. Although the new proof is introduced to get rid of the constant factor 222 in front of $\left\lVert X\beta-X\beta_{I_{0}}\right\rVert_{2}/\sqrt{n}$ in (22) in [48], we are not optimizing the constants in this paper, cf. the proof of the original Theorem 5.1 in [48]. For example, the constant 333 in (40) can be further reduced following [11] using (126). One can also use $\textnormal{{RE}}(s_{0},6,X)$ in Case 2 to further tighten the constants in (52) and (56) for Case 3, for example, by considering in Case 2,
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}$$\leq$ $3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1},$
instead of (137); and in Case 3, instead of (51),
@@
Now clearly $\frac{\Lambda_{\max}(s-s_{0})}{\Lambda_{\min}(2s_{0})}\geq\frac{\sqrt{\Lambda_{\max}(s-s_{0})}}{\sqrt{\Lambda_{\min}(2s_{0})}}$ for $s\geq 2s_{0}$ since
$\Lambda_{\min}(2s_{0})\leq\Lambda_{\min}(s_{0})\leq\Lambda_{\max}(s_{0})\leq\Lambda_{\max}(s-s_{0}),$
and $1/\sqrt{\Lambda_{\min}(2s_{0})}\leq\sqrt{2}K(s_{0},1)$. Compared to Case 1 and Case 2, where $\textnormal{{RE}}(s_{0},k_{0},X)$ holds for $k_{0}=3,4$ respectively, we have an extra term on the bound for $\left\lVert h_{T_{01}}\right\rVert_{2}$ in (56), namely, $\ell(s_{0})\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}$ that is bounded as follows: denote by $D_{1,c}=3\Lambda_{\max}(s-s_{0})/d_{0}$,
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}$$\leq$ $\frac{3\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}}{\lambda_{n}\sqrt{s_{0}}}\leq D_{1,c}\lambda\sigma\sqrt{s_{0}}$
[…]
@@
Under the UUP condition as in Definition I.1:
$\forall s_{0}\leq s,\quad\theta_{s_{0},2s_{0}}+\delta_{2s_{0}}\leq\theta_{s,2s}+\delta_{2s}<1$
since $\theta_{s,2s}$ and $\delta_{s}$ are nondecreasing in $s$. Thus we have
$\theta_{s_{0},2s_{0}}<1-\delta_{2s_{0}}\leq\Lambda_{\min}(2s_{0}).$ (57)
When such tight bounds are not available, we can still use the bounds in (29) to control $\theta_{s_{0},2s_{0}}$. Moreover, suppose $\theta_{s_{0},2s_{0}}<\Lambda_{\min}(2s_{0})$, which holds under UUP (57), one can get rid of the factor $\sqrt{\Lambda_{\max}(s_{0})}$ in (54), by bounding $\left\lVert h_{T_{01}}\right\rVert_{2}$ for Case 3 following Lemma 3.1 instead. The proof of Lemma 3.1 is included in the supplementary Section F for self-containment.
@@ 4 On Type II errors and $\ell_{2}$-loss optimality
Let $A_{0}:=\{j:|\beta_{j}|>\sqrt{2\log p/n}\sigma\}$. Let $T_{0}$ denote positions of the $s_{0}$ largest coefficients of $\beta$ in absolute values, where $s_{0}$ is defined in (13). Let $a_{0}=\left|A_{0}\right|$ denote the cardinality of $A_{0}$ (see also (15)). Then $\forall c^{\prime}>1/2$, we have |$\left|\{j\in T_{0}^{c}:|\beta_{j}|\geq\sqrt{\log p/(c^{\prime}n)}\sigma\}\right|\leq(2c^{\prime}-1)(s_{0}-a_{0})$.
@@
|$\left|\{j\in T_{0}^{c}:|\beta_{j}|\geq\sigma\sqrt{2/(n)}\}\right|\leq(\log p-1)(s_{0}-a_{0})$ (58)
We first show in Lemma 4.2 that under no restriction on $\beta_{\min}$, we achieve an oracle bound on the $\ell_{2}$ loss, which depends only on the $\ell_{2}$ loss of the initial estimator on the set $T_{0}$. Bounds in Lemma 2.3 is a special case of (59). We prove Lemma 4.2 in Section B.2. In Lemma 4.3, we impose a lower bound on $\beta_{\min,A_{0}}$ as in (61) in order to recover the subset of variables in $A_{0}$, while achieving the nearly ideal $\ell_{2}$ loss with a sparse model $I$.
@@
(A deterministic result on the bias.) Let $\beta_{\text{\rm init}}$ be an initial estimator. Let $h=\beta_{\text{\rm init}}-\beta_{T_{0}}$ be as defined in Lemma 2.3. Let $\lambda:=\sqrt{2\log p/n}.$ Suppose we choose a thresholding parameter $t_{0}$ and set
@@
(Oracle Ideal MSE with $\ell_{\infty}$ bounds) Suppose (6) and (5) hold. Let $\beta_{\text{\rm init}}$ be an initial estimator. Let $h=\beta_{\text{\rm init}}-\beta_{T_{0}}$ and $\lambda:=\sqrt{2\log p/n}$. Suppose on some event $Q_{c}$, for $\beta_{\min,A_{0}}$ as defined in (15), it holds that
@@
Now we choose a thresholding parameter $t_{0}$ such that on $Q_{c}$,
@@
holds and set $I=\{j:\left|\beta_{j,\text{\rm init}}\right|\geq t_{0}\}$. Then we have on ${\mathcal{T}}_{a}\cap Q_{c}$,
@@
Let $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$. Then for $\breve{s}_{0}\leq s$, we have on ${\mathcal{T}}_{a}\cap Q_{c}$,
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}\leq\left(1+\frac{(\Lambda_{\max}(2s)-\Lambda_{\min}(2s))^{2}+8(1+a)}{2\Lambda_{\min}^{2}(|I|)}\right)\breve{s}_{0}\lambda^{2}\sigma^{2},$ (65)
where $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$ is the 0-extended version of $\hat{\beta}_{I}$ such that $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}=0$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}=\hat{\beta}_{I}$.
In Theorem 4.4, we show that one can indeed recover a subset $A_{0}$ of variables accurately, for $A_{0}$ as defined in Proposition 4.1, when $\beta_{\min,A_{0}}:=\min_{j\in A_{0}}|\beta_{j}|$ is large enough (relative to the $\ell_{2}$ loss of an initial estimator $\beta_{\text{\rm init}}$ under the RE condition on the set $A_{0}$); in addition, a small number of extra variables from $T_{1}\subset T_{0}^{c}:=[p]\setminus T_{0}$ are also possibly included in the model $I$. We mention in passing that changing the coefficients of $\beta_{A_{0}}$ will not change the values of $s_{0}$ or $a_{0}$, so long as their absolute values stay strictly above $\lambda\sigma$.
Given all other parameters held invariant, a larger $\beta_{\min,A_{0}}$ (e.g., a larger $C_{a}$ for $\beta$ in Figure 1) will result in a tighter bound on the bias $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}$ in view of (60). When $\beta_{\min,A_{0}}^{2}$ dominates the upper bound on the RHS of (60), then $A_{0}\cap I^{c}=\emptyset$, since the loss of a single variable from $A_{0}$ will already saturate this upper bound in the order of $O(\lambda^{2}\sigma^{2}s_{0})$. Hence when $\beta_{\min,A_{0}}$ is strong enough in the sense of (4.3), there will be no false negatives from the set $A_{0}$, leading to the stronger and tighter bounds in (63) and (64) that also control false positives and the bias. This is validated in our experiments in Section 5.
Theorem 4.4 is an immediate corollary of Theorem 2.1 and Lemma 4.3, except that we now let $\breve{s}_{0}=s_{0}$ everywhere and assume having an upper estimate $\breve{D}_{1}$ (resp. $\breve{D}_{0}$) of $D_{1}$ (resp. $D_{0}$), so as not to depend on an “oracle” telling us an exact value. We provide a direct proof here.
Theorem 4.4.
Suppose $\textnormal{{RE}}(s_{0},4,X)$ holds. Suppose (6) and (5) hold, and $\theta_{s_{0},2s_{0}}<\infty$. Choose $\lambda_{n}\geq b\lambda_{\sigma,a,p}$, where $b\geq 2$. Let $D_{0},D_{1}$ be as in (2.1) and (26). Let $T_{0}$ denote the largest $s_{0}$ coordinates of $\beta$ in absolute values. Let $\beta_{\text{\rm init}}$ be the Lasso estimator as in (2). Let $\beta_{\min,A_{0}}:=\min_{j\in A_{0}}|\beta_{j}|$. Suppose for some constants $\breve{D}_{1}>D_{1}$, $\breve{D}_{0}>D_{0}$ and $\lambda:=\sqrt{2\log p/n}$,
$\beta_{\min,A_{0}}\geq D_{0}\lambda\sigma\sqrt{s_{0}}+(\breve{D}_{1}\wedge\breve{D}_{0})\lambda\sigma.$ (66)
Let $T_{1}$ be the largest $s_{0}$ positions of $\beta_{\text{\rm init}}$ outside of $T_{0}$. Let $T_{01}=T_{0}\cup T_{1}$. Choose a thresholding parameter $t_{0}$ and set
Then on event ${\mathcal{T}}_{a}$,
$A_{0}\subset I,\quad\;I\cap T_{01}^{c}=\emptyset,\;\text{and}\;\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}\leq\lambda\sigma\sqrt{s_{0}-a_{0}},$
and (65) holds with $\breve{s}_{0}=s_{0}$ and $\left\lvert I\right\rvert\leq 2s_{0}$.
Proof.
By Lemma 3.4, we have for $h=\beta_{\text{\rm init}}-\beta_{0}$, where $\beta_{0}=\beta_{T_{0}}$, on event ${\mathcal{T}}_{a}$,
$\beta_{\min,A_{0}}-\left\lVert h_{A_{0}}\right\rVert_{\infty}$$\geq$ $t_{0}\text{ or the stronger }\beta_{\min,A_{0}}\geq\left\lVert h_{A_{0}}\right\rVert_{2}+t_{0};$
$\text{Then }\;\forall j\in A_{0},\;\left\lvert\beta_{\text{\rm init},j}\right\rvert$$\geq$ $\left\lvert\beta_{\min,A_{0}}-\left\lVert h_{A_{0}}\right\rVert_{\infty}\right\rvert>t_{0},$ (68)
@@
we can then eliminate all variables in $T_{01}^{c}$ from model $I$, where $\beta_{\text{\rm init},T_{0}^{c}}=h_{T_{0}^{c}}$ and $T_{1}\subset T_{0}^{c}$. To see the second inequality in (69), recall the $k$th largest value of $\left\lvert h_{T_{0}^{c}}\right\rvert$ obeys $\left|h_{T_{0}^{c}}\right|_{(k)}\leq\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/k$. Hence by definition of $T_{1}$ and (67),
$\left\lVert\beta_{\text{\rm init},T_{01}^{c}}\right\rVert_{\infty}$ $=$ $\left|h_{T_{0}^{c}}\right|_{(s_{0}+1)}\leq\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/(s_{0}+1)<D_{1}\lambda\sigma;$
@@
$\beta_{\min,A_{0}}$$\geq$ $\left\lVert h_{T_{0}}\right\rVert_{2}+t_{0}>\left\lVert h_{T_{0}}\right\rVert_{\infty}+(D_{0}\wedge D_{1})\lambda\sigma;$ (70)
Then none of the variables in $T_{01}^{c}$ can be chosen in the thresholding step and hence $I\cap T_{01}^{c}=\emptyset$. The rest follows from the proof of Lemma 4.3. ∎∎

### v3

@@ 3 Proof sketch for the main result
Under the settings of Theorem 2.4, we have under event ${\mathcal{T}}_{a}$,
$\text{\bf Case 1}:$ $\left\lVert h\right\rVert_{2}\leq 2\left\lVert h_{T_{01}}\right\rVert_{2}\leq 2D_{0,a}\lambda\sigma\sqrt{s_{0}}\;\text{where}$
$D_{0,a}$ $\leq$ $\sqrt{2}\big(\sqrt{\Lambda_{\max}(s-s_{0})}K(s_{0},3)+3d_{0}K^{2}(s_{0},3)\big);$ (52)
$\text{\bf Case 2}:$ $\left\lVert h\right\rVert_{2}\leq\sqrt{5}\left\lVert h_{T_{01}}\right\rVert_{2}\leq\sqrt{10}K^{2}(s_{0},4)d_{0}\lambda\sigma\sqrt{s_{0}};$
$\text{\bf Case 3}:$ $\left\lVert h_{T_{01}}\right\rVert_{2}\leq\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\big(\left\lVert Xh\right\rVert_{n}+\frac{\ell(s_{0})}{\sqrt{s_{0}}}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\big)\leq D\lambda\sigma\sqrt{s_{0}},$ (53)
$\text{where}\;D:=\sqrt{\frac{\Lambda_{\max}(s-s_{0})}{\Lambda_{\min}(2s_{0})}}\big(1+\frac{3\ell(s_{0})}{d_{0}}\sqrt{\Lambda_{\max}(s-s_{0})}\big)$
and $\ell(s_{0})$ is as defined in (D.2). Moreover, under event ${\mathcal{T}}_{a}$, we have for $D_{0}$ in (23) and $D_{1}$ in (27): $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq D_{1}\lambda\sigma s_{0}$,
@@
[…]
$\left\lVert\beta_{\text{\rm init}}-\beta\right\rVert_{2}$ $\leq$ $\left\lVert h\right\rVert_{2}+\lambda\sigma\sqrt{s_{0}}\leq[\sqrt{D_{0}^{2}+D_{1}^{2}}+1]\lambda\sigma\sqrt{s_{0}}.$
@@
One can also use $\textnormal{{RE}}(s_{0},6,X)$ in Case 2 to further tighten the constants in (51) and (53) for Case 3, for example, by considering
$\text{ in Case 2},\quad\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\leq 3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1},$
instead of (136), and in Case 3, instead of (50),
@@
Now clearly $\frac{\Lambda_{\max}(s-s_{0})}{\Lambda_{\min}(2s_{0})}\geq(\frac{\Lambda_{\max}(s-s_{0})}{\Lambda_{\min}(2s_{0})})^{1/2}$ for $s\geq 2s_{0}$ since
$\Lambda_{\min}(2s_{0})\leq\Lambda_{\min}(s_{0})\leq\Lambda_{\max}(s-s_{0}),\;\text{ and }1/\sqrt{\Lambda_{\min}(2s_{0})}\leq\sqrt{2}K(s_{0},1).$
Compared with Case 1 and Case 2, where $\textnormal{{RE}}(s_{0},k_{0},X)$ holds for $k_{0}=3,4$ respectively, we have an extra term on the bound for $\left\lVert h_{T_{01}}\right\rVert_{2}$ in (53), namely, $\ell(s_{0})\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}\asymp\lambda\sigma\sqrt{s_{0}}$; cf. (51) and (D.2).
@@
Under the UUP condition as in Definition G.1:
$\forall s_{0}\leq s,\quad\theta_{s_{0},2s_{0}}+\delta_{2s_{0}}\leq\theta_{s,2s}+\delta_{2s}<1$ (54)
since $\theta_{s,2s}$ and $\delta_{s}$ are nondecreasing in $s$. Thus we have $\theta_{s_{0},2s_{0}}<1-\delta_{2s_{0}}\leq\Lambda_{\min}(2s_{0})$. When such tight bounds are not available, we can still use the bounds in (30) to control $\theta_{s_{0},2s_{0}}$. Moreover, suppose $\theta_{s_{0},2s_{0}}<\Lambda_{\min}(2s_{0})$, which holds under UUP (54), one can get rid of the factor $\sqrt{\Lambda_{\max}(s_{0})}$ in (D.2), where recall $\ell(s_{0})=({\theta_{s_{0},2s_{0}}}/{\sqrt{\Lambda_{\min}(2s_{0})}})\wedge\sqrt{\Lambda_{\max}(s_{0})}$, by bounding $\left\lVert h_{T_{01}}\right\rVert_{2}$ for Case 3 following (38) instead of the supplementary Lemma D.2.
@@ 4 On Type II errors and $\ell_{2}$-loss optimality
Let $A_{0}:=\{j:|\beta_{j}|>\sqrt{2\log p/n}\sigma\}$. Let $T_{0}$ denote positions of the $s_{0}$ largest coefficients of $\beta$ in absolute values, where $s_{0}$ is defined in (13). Let $a_{0}=\left|A_{0}\right|$ denote the cardinality of $A_{0}$ (see also (15)). Then $\forall c^{\prime}>1/2$, we have $\left|\{j\in T_{0}^{c}:|\beta_{j}|\geq\sqrt{\log p/(c^{\prime}n)}\sigma\}\right|\leq(2c^{\prime}-1)(s_{0}-a_{0})$.
@@
$\left|\{j\in T_{0}^{c}:|\beta_{j}|\geq\sigma\sqrt{2/(n)}\}\right|\leq(\log p-1)(s_{0}-a_{0}).$ (55)
We prove Proposition 4.1 in Section 6.3. We first show in Lemma 4.2 that under no restriction on $\beta_{\min}$, we achieve an oracle bound on the $\ell_{2}$ loss, which depends only on the $\ell_{2}$ loss of the initial estimator on the set $T_{0}$. Bounds in Lemma 2.3 are special cases of (56). In Lemma 4.3, we impose a lower bound on $\beta_{\min,A_{0}}$ as in (58) in order to recover the subset of variables in $A_{0}$, while achieving the nearly ideal $\ell_{2}$ loss with a sparse model $I$ in view of Lemma 2.7. We prove Lemma 4.2 and 4.3 in Sections 6.6 and 6.7 respectively.
@@
(A deterministic result on the bias.) Let $\beta_{\text{\rm init}}$ be an initial estimator. Let $\lambda:=\sqrt{2\log p/n}$ and $h\in\mathbb{R}^{p}$ be as defined in Lemma 2.3. Suppose we choose a thresholding parameter $t_{0}$ and set
@@
(A deterministic result on the bias and model selection.) Let $\beta_{\text{\rm init}}$ be an initial estimator. Let $\lambda:=\sqrt{2\log p/n}$ and $h\in\mathbb{R}^{p}$ be as defined in Lemma 2.3. Suppose for $\beta_{\min,A_{0}}$ as defined in (15), it holds that
@@
Now we choose a thresholding parameter $t_{0}$ such that
@@
holds and set $I=\{j:\left|\beta_{j,\text{\rm init}}\right|\geq t_{0}\}$. Then we have
$A_{0}\subset I$ $\text{ and }\quad|I\cap T_{0}^{c}|\leq\breve{s}_{0},\quad\text{ and hence}$ (60)
$|I|\leq s_{0}+\breve{s}_{0}$ $\text{ and }\quad\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\leq(s_{0}-a_{0})\lambda^{2}\sigma^{2}.$ (61)
Choosing the set $A_{0}$ as in Proposition 4.1 is rather arbitrary; one could for example, consider the set of variables that are strictly above $\lambda\sigma/2$. In Theorem 4.4, we show that one can indeed recover a subset $A_{0}$ of variables accurately, for $A_{0}$ as defined in Proposition 4.1, when $\beta_{\min,A_{0}}:=\min_{j\in A_{0}}|\beta_{j}|$ is large enough (relative to the $\ell_{2}$ loss of an initial estimator $\beta_{\text{\rm init}}$ under the RE condition on the set $A_{0}$); in addition, a small number of extra variables from $T_{1}\subset T_{0}^{c}:=[p]\setminus T_{0}$ are also possibly included in the model $I$. We mention in passing that changing the coefficients of $\beta_{A_{0}}$ will not change the values of $s_{0}$ or $a_{0}$, so long as their absolute values stay strictly above $\lambda\sigma$; See Section 4.2. Compared with the almost exact sparse recovery result in Theorem 1.1 [47], we have relaxed the restriction on $\beta_{\min}$: rather than requiring all non-zero entries to be large in absolute values, we only require those in a subset $A_{0}$ to be recovered to be large.
Theorem 4.4.
Under the settings of Theorem 2.1, let $\beta_{\text{\rm init}}$ be the Lasso estimator as in (2). Let $T_{0}$ denote the largest $s_{0}$ coordinates of $\beta$ in absolute values. Let $T_{1}$ be the largest $s_{0}$ positions of $\beta_{\text{\rm init}}$ outside of $T_{0}$. Let $T_{01}=T_{0}\cup T_{1}$. Suppose (5), (6), and $\textnormal{{RE}}(s_{0},4,X)$ condition hold. Let $D_{0},D_{1}$ be as in (23) and (27). Suppose for some constants $\breve{D}_{1}>D_{1}$, $\breve{D}_{0}>D_{0}$,
$\beta_{\min,A_{0}}:=\min_{j\in A_{0}}|\beta_{j}|\geq D_{0}\lambda\sigma\sqrt{s_{0}}+(\breve{D}_{1}\wedge\breve{D}_{0})\lambda\sigma$ (62)
where $\lambda:=\sqrt{2\log p/n}$. Choose a thresholding parameter $t_{0}$ and set
Let $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$ be the 0-extended version of $\hat{\beta}_{I}$ such that $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}=0$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}=\hat{\beta}_{I}$. Then on event ${\mathcal{T}}_{a}$, we have $\left\lvert I\right\rvert\leq 2s_{0}$,
$A_{0}\subset I,\quad\;I\cap T_{01}^{c}=\emptyset,\;\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}\leq\lambda\sigma\sqrt{s_{0}-a_{0}},\text{ and}$
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}\leq\big(1+\frac{(\Lambda_{\max}(2s)-\Lambda_{\min}(2s))^{2}+8(1+a)}{2\Lambda_{\min}^{2}(|I|)}\big)s_{0}\lambda^{2}\sigma^{2}.$ (63)
Proof.
By definition of $T_{0}$, we have $T_{0}\subset S$ and recall ${S_{{\mathcal{D}}}}={\mathcal{D}}\cap S$. Moreover, $I\cap{S_{{\mathcal{D}}}}=\emptyset$ by definition. By Lemma 3.3, we have for on ${\mathcal{T}}_{a}$,
@@
we can then eliminate all variables in $T_{01}^{c}$ from model $I$, where $\beta_{\text{\rm init},T_{0}^{c}}=h_{T_{0}^{c}}$ and $T_{1}\subset T_{0}^{c}$. To see the second inequality in (66), recall the $k$th largest value of $\left\lvert h_{T_{0}^{c}}\right\rvert$ obeys $\left|h_{T_{0}^{c}}\right|_{(k)}\leq\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/k$ for $k\geq 1$. Hence by definition of $T_{1}$,
[…]
@@
$\beta_{\min,A_{0}}$ $\geq$ $\left\lVert h_{T_{0}}\right\rVert_{2}+t_{0}>\left\lVert h_{T_{0}}\right\rVert_{\infty}+(D_{0}\wedge D_{1})\lambda\sigma.$ (67)
Then none of the variables in $T_{01}^{c}$ can be chosen in the thresholding step and hence $I\cap T_{01}^{c}=\emptyset$. Hence, $\left|I\right|\leq 2s_{0}$. On event ${\mathcal{T}}_{a}$, (33) holds by Lemma 2.7. Moreover, $\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}$ is bounded as in Lemma 2.8 given $|I|+|{S_{{\mathcal{D}}}}|=|I\cup{S_{{\mathcal{D}}}}|\leq\left\lvert S\right\rvert+\left\lvert I\cap S^{c}\right\rvert\leq s+|I\cap T_{0}^{c}|\leq s+s_{0}\leq 2s$. Thus (63) follows immediately. ∎∎